### v1

# Does Machine Unlearning Truly Remove Model Knowledge? A Framework for Auditing Unlearning in LLMs

Abstract
In recent years, Large Language Models (LLMs) have achieved remarkable advancements, drawing significant attention from the research community. Their capabilities are largely attributed to large-scale architectures, which require extensive training on massive datasets. However, such datasets often contain sensitive or copyrighted content sourced from the public internet, raising concerns about data privacy and ownership. Regulatory frameworks, such as the General Data Protection Regulation (GDPR), grant individuals the right to request the removal of such sensitive information. This has motivated the development of machine unlearning algorithms that aim to remove specific knowledge from models without the need for costly retraining. Despite these advancements, evaluating the efficacy of unlearning algorithms remains a challenge due to the inherent complexity and generative nature of LLMs. In this work, we introduce a comprehensive auditing framework for unlearning evaluation, comprising three benchmark datasets, six unlearning algorithms, and five prompt-based auditing methods. By using various auditing algorithms, we evaluate the effectiveness and robustness of different unlearning strategies. To explore alternatives beyond prompt-based auditing, we propose a novel auditing technique based on intermediate activation perturbation. The complete framework will be open-sourced upon manuscript acceptance.
Does Machine Unlearning Truly Remove Model Knowledge? A Framework for Auditing Unlearning in LLMs
Haokun Chen1,2,3, Yueqi Zhang4, Yuan Bi4, Yao Zhang1,3, Tong Liu1,3, Jinhe Bi1,3, Jian Lan1,3, Jindong Gu5, Claudia Grosser2,4, Denis Krompass2, Nassir Navab4, Volker Tresp1,3 1 LMU Munich, Germany 2 Siemens AG, Munich, Germany 3 Munich Center for Machine Learning (MCML), Germany 4 Technical University of Munich, Germany 5 University of Oxford, UK
1 Introduction
Large language models (LLMs) have seen rapid advancements recently, resulting in improved performance and widespread adoption across numerous applications. These advancements are largely attributed to their large-scale architectures, which require training on datasets containing billions of tokens Kaplan et al. (2020). These datasets are typically constructed from large-scale corpora of publicly available internet text. However, such corpora often inadvertently include personally identifiable information (PII) or copyrighted material, which are considered sensitive and generally unsuitable for commercial use due to legal and ethical constraints. To comply with local regulations (e.g., GDPR li (2)) and internal organizational policies, it is often necessary to remove sensitive information from trained models.
Machine unlearning has emerged as a promising solution to this problem Cao and Yang (2015); Bourtoule et al. (2021a). This work is motivated by the legal framework proposed by the European Union, namely the GDPR li (2), which grants individuals the right to request the removal of their personal data in trained models. In particular, approximate unlearning seeks to remove specific knowledge from a model without the need for retraining from scratch Zhang et al. (2024a); Eldan and Russinovich (2023a); Izzo et al. (2021), while ensuring that the resulting model closely approximates a retrained counterpart within a bounded error. This approach is especially appealing in the context of LLMs, where full retraining is prohibitively expensive. Despite the development of numerous unlearning algorithms, few studies have systematically assessed their effectiveness and robustness Qi et al. (2024). Recent research has shown that many of these methods can be easily circumvented using simple paraphrasing attacks Shi et al. (2024).
To advance research on evaluating existing unlearning algorithms, we introduce a comprehensive framework for auditing unlearning in LLMs. The proposed framework incorporates 3 benchmark datasets, 6 representative unlearning algorithms, and 5 prompt-based auditing strategies. Leveraging this setup, we perform an extensive evaluation of the effectiveness and robustness of various unlearning methods. To explore alternatives to prompt-based auditing, we introduce a novel technique that perturbs intermediate model outputs to detect residual traces of forgotten information. Our key contributions are as follows:
  • We propose an prompt-based auditing framework for evaluating unlearning in LLMs.
  • We propose a novel activation perturbation-based auditing method to detect memorized traces of removed content.
  • We conduct extensive experiments within our framework and provide an in-depth analysis of the effectiveness and limitations of current unlearning algorithms for LLMs.
2 Related Work
2.1 Machine Unlearning in LLMs
Machine unlearning has garnered significant attention in the context of LLMs. Various approaches for targeted knowledge removal have been proposed:  Eldan and Russinovich (2023a) removed Harry Potter-related knowledge by finetuning LLMs on corpora with replaced keywords.  Zhang et al. (2024b) proposed to steer model preferences in the negative direction to reduce memorization.  Wang et al. (2024) used reversed knowledge distillation to eliminate personal information.  Feng et al. (2024) introduced a reweighted gradient ascent method for unlearning, and  Pawelczyk et al. (2023) utilized in-context unlearning examples.  Liu et al. (2024a); Bhaila et al. (2024) adapted input embeddings associated with the unlearning target, while Li et al. (2024); Tamirisa et al. (2024); Huu-Tien et al. (2024); Ashuach et al. (2024) proposed interventions in the model activation space. From a safety perspective, several studies have targeted the unlearning of harmful or undesirable model behaviors Zhang et al. (2024c); Yao et al. (2023); Liu et al. (2024c), while others have addressed backdoor unlearning Liang et al. (2024); Liu et al. (2024b) to enhance model integrity.
2.2 Evaluation of LLM Unlearning
While advances have been made in developing machine unlearning algorithms for LLMs, rigorous methodologies for auditing the efficacy of unlearning remain understudied. Recent work has demonstrated that unlearned knowledge can be recovered using techniques such as adversarial soft token attacks Schwinn et al. (2024) or few-shot in-context prompting Doshi and Stickland (2024). When model parameters are accessible, recovery can also be achieved through model quantization Zhang et al. (2024d) or retraining on partially unlearned datasets Łucki et al. (2024); Hu et al. (2024).  Lynch et al. (2024) compared eight unlearning evaluation methods and found persistent similarity in the latent representations of unlearned models.  Shi et al. (2024) conducted a comprehensive analysis across six evaluation dimensions using news and book datasets, highlighting the fragility of unlearning to paraphrasing and unrelated data.  Qi et al. (2024) showed that fine-tuning on unrelated data could inadvertently restore previously removed knowledge. Several benchmarks have been developed to evaluate the existing unlearning algorithms.  Maini et al. (2024) proposed a benchmark centered on fictitious author profiles, while Jin et al. (2024) introduced the Real-World Knowledge Unlearning (RWKU) benchmark, which targets knowledge removal of 200 well-known individuals. Complementary to this, the WDMP benchmark Li et al. (2024) focuses on the unlearning of sensitive knowledge related to biosecurity and cybersecurity threats.
Figure 1: The proposed activation perturbation-based algorithm (ActPert) for auditing unlearning in LLMs.
3 Preliminaries
Large Language Models (LLMs) operate within an auto-regressive framework, generating text by predicting the next token in a sequence based on preceding tokens. Given an input sequence of tokens $x_{1:t}$, the model computes the conditional probability distribution $p(x_{t+1}|x_{1:t})$ over the vocabulary at each time step. The overall likelihood of the sequence is given by:
$\log p(x_{t+1}|x_{1:t})=\sum_{t=1}^{T}\log p(x_{t}|x_{1:t-1})$ (1)
During inference, tokens are generated iteratively: at each step, the model samples the next token $x_{t+1}$ from the predicted distribution (using methods such as greedy decoding or nucleus sampling Holtzman et al. (2019)), then appends it to the existing context $x_{1:t}$ for the subsequent prediction.
Machine unlearning refers to the process of removing the influence of specific data from a trained model. Consider a machine learning model $f$ trained on a dataset $D_{train}$. When a data owner requests the removal of a subset $D_{u}\in D_{train}$, the goal of machine unlearning is to produce a modified model $f_{u}$ that behaves as if it had never been trained on $D_{u}$. Unlearning techniques generally fall into two categories: exact unlearning, which seeks to fully eliminate the impact of the forgotten data, and approximate unlearning, which aims for partial or probabilistic removal. We introduce them in the following.
Exact unlearning guarantees that the output distribution of the unlearned model $f_{u}$ is statistically indistinguishable from that of a reference model $f_{ref}$, which is trained from scratch on the retained dataset $D_{r}=D_{train}-D_{u}$. This strong guarantee of data removal is formalized as:
$p(f_{u}(x)=y)$ $=p(f_{ref}(x)=y)\quad$ (2)
$s.t.\quad\forall(x,y)\in D_{train}.$
While retraining from scratch is the most direct method to achieve exact unlearning, it is often computationally infeasible for large-scale models such as LLMs. One solution is to partition $D_{train}$ into overlapping subsets and train an ensemble of models Bourtoule et al. (2021b). Upon receiving an unlearning request, only the models that contain the targeted records need to be retrained.
Approximate unlearning relaxes the requirement of strict distributional equivalence. It seeks to ensure that the behavior of $f_{u}$ closely approximates $f_{ref}$ within a tolerable margin of error, often quantified through empirical metrics or probabilistic bounds.
In the context of LLMs, approximate unlearning is typically realized through information overwriting Eldan and Russinovich (2023a); Wang et al. (2024), behavioral steering Feng et al. (2024), or model editing—via weight or activation modifications Liu et al. (2024a); Bhaila et al. (2024); Li et al. (2024); Tamirisa et al. (2024); Huu-Tien et al. (2024); Ashuach et al. (2024); Meng et al. (2022a, b). These methods aim to diminish or redirect the model’s reliance on the forgotten data without necessitating a full retraining cycle.
4 Proposed Method
Before introducing the proposed framework for unlearning auditing, we introduce Activation Perturbation-based Auditing (ActPert), a method for probing residual knowledge in unlearned language models. A schematic overview of the proposed approach is shown in Figure 1. Our method builds on recent advances in activation engineering for LLMs Arditi et al. (2024), which compute “refusal directions” by contrasting activations between harmful and harmless prompts to reduce a model’s tendency to refuse answering.
Analogously, we treat unlearning targets as sensitive queries and seek to perturb their input representations such that they become effectively insensitive, thereby increasing the chances of eliciting meaningful responses. Concretely, we inject random noise into the token embeddings corresponding to the unlearning target (e.g., the phrase Harry Potter in the prompt "Who are Harry Potter’s two best friends?"). This noise injection prevents the model from directly attending to the sensitive content during inference, resulting in a set of $n_{p}$ perturbed versions of the original query in the embedding space.
For each transformer layer $k$ in the unlearned LLM, we compute an activation perturbation $\delta_{k}$ as the difference between the layer activation of the original (unperturbed) query, denoted $A_{k}$, and the mean activation across the perturbed variants, denoted $\hat{A}_{k}^{i}$:
$\delta_{k}=A_{k}-\frac{1}{n_{p}}\sum_{n_{p}}\hat{A}_{k}^{i}$ (3)
These layer-wise perturbations $\delta_{k}$ are then reintroduced into the model during autoregressive generation. By modifying the model’s internal activations at inference time, this intervention allows us to assess whether residual knowledge of the unlearning target still influences the model’s outputs.
5 Experiments and Analyses
In this section, we provide details about the proposed unlearning auditing framework, which encompasses commonly used unlearning algorithms as well as established benchmarks. We begin by outlining representative benchmarks, unlearning methods, and auditing methods used in our evaluation. Following this, we report validation results from multiple auditing algorithms, including our proposed ActPert, and provide a comparative analysis to assess their effectiveness in detecting residual knowledge.
5.1 Unlearning Benchmarks
In this section, we introduce the unlearning benchmarks included in our framework:
  • WHP Eldan and Russinovich (2023b). We audit a pretrained language model that was finetuned to unlearn knowledge related to Harry Potter. As the authors only released the unlearned model checkpoint without the original unlearning dataset $D_{u}$, we use GPT-4o to generate Q&A pairs for auditing. Specifically, we prompt GPT-4o to produce factual, short-form questions about Harry Potter. These generated questions are then submitted to both the pretrained and the unlearned models. Questions for which both models produce either correct or incorrect answers are filtered out. Further details about the filtering process are provided in the appendix. In total, we select 35 Q&A pairs for auditing.
  • TOFU Maini et al. (2024). TOFU is an unlearning benchmark constructed from autobiographies of 200 fictitious authors generated using GPT-4o, ensuring no prior presence in LLM pretraining. The pretrained language model is first finetuned on the full dataset, then unlearning is applied to either 1% (2 authors) or 5% (10 authors) of the training data. We follow the original paper’s hyperparameter settings for both fine-tuning and unlearning. While the original TOFU dataset consists of long-form Q&A pairs, we adopt GPT-4o to generate short-form questions and maintain consistency with other benchmarks. We filter out questions that the finetuned model answers incorrectly. This process yields 16 Q&A pairs for auditing the model unlearned with 1% of TOFU and 80 Q&A pairs for the model unlearned with 5%.
  • RWKU Jin et al. (2024). RWKU is a benchmark focused on real-world knowledge unlearning. It constructs Q&A pairs about well-known public figures as unlearning targets. This approach avoids the need for retraining, as such knowledge is prevalent in open-source LLMs. The unlearning dataset $D_{u}$ consists of biographical texts about these individuals, and RWKU includes Q&A pairs for evaluating different levels of unlearning effectiveness. To examine the impact of unlearning dataset size, we select 10, 20, and 30 individuals as unlearning targets and evaluate all unlearned models using the same 10 individuals.
5.2 Model Architecture
For the WHP benchmark, we use the model checkpoints provided by the original authors, which is based on the Llama-2-Chat111meta-llama/Llama-2-7b-chat-hf architecture. For the TOFU benchmark, we adopt the same Llama-2-Chat model as the base model and finetune it on the full TOFU training set. For the RWKU benchmark, we perform unlearning on both the pretrained Llama-3-Instruct222meta-llama/Meta-Llama-3-8B-Instruct and Phi-3-mini-instruct333microsoft/Phi-3-mini-4k-instruct models. All model checkpoints are obtained from open-sourced HuggingFace library.
5.3 Unlearning Algorithms
In this section, we describe the unlearning algorithms that are evaluated in the framework:
  • Gradient Ascent (GA) minimizes the probability that the target model $f_{u}$ makes correct predictions on the unlearning set $D_{u}$.
    $\mathcal{L}_{GA}=\frac{1}{|D_{u}|}\sum_{x\in D_{u}}\ell(x,f_{u}).$ (4)
  • Gradient Difference (GD) is a variant of GA that incorporates an additional loss term to preserve performance on the retain set $D_{r}$.
    $\mathcal{L}_{GD}=-\frac{1}{|D_{r}|}\sum_{x\in D_{r}}\ell(x,f_{u})+\mathcal{L}_{GA}(D_{u},f_{u}).$ (5)
  • Knolwedge Distillation (KD) extends GA by minimizing the KL divergence between the output token probabilities of the fine-tuned model ($f_{ft}$) and the unlearned model on the retained dataset $D_{r}$.
    $\begin{split}\mathcal{L}_{KD}=&-\frac{1}{|D_{r}|}\sum_{x\in D_{r}}\mathcal{L}_{KL}(f_{ft}(x)||f_{u}(x))\\ &+\mathcal{L}_{GA}(D_{u},f_{u}).\end{split}$ (6)
  • Rejection Tuning (RT) Maini et al. (2024) aligns the model to refuse when queried about target knowledge. This is achieved by constructing $D_{u}^{\text{idk}}$, where the responses to questions in the unlearning set $D_{u}$ are replaced with I don’t know or similar refusal-style responses.
    $\begin{split}\mathcal{L}_{RT}=&-\frac{1}{|D_{r}|}\sum_{x\in D_{r}}\ell(x,f_{u})\\ &-\frac{1}{|D_{u}^{idk}|}\sum_{x\in D_{u}^{idk}}\ell(x,f_{u}).\end{split}$ (7)
  • Direct Preference Optimization (DPO) (Rafailov et al., 2023) aligns the model to avoid providing accurate target knowledge. Specifically, counterfactual responses are used as positive examples $y_{w}$, consisting of model-generated fabrications about the unlearning target, while negative responses $y_{l}$ are drawn based on the ground-truth knowledge:
    $\begin{split}\mathcal{L}_{DPO}=&-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}_{u}}[\log\sigma\\ &(\beta\log\frac{f_{u}(y_{w}\mid x)}{f_{ft}(y_{w}\mid x)}-\beta\log\frac{f_{u}(y_{l}\mid x)}{f_{ft}(y_{l}\mid x)})]\end{split}$ (8)
  • Negative Preference Optimization (NPO) (Zhang et al., 2024b) is a variant of DPO that retains only the negative examples, i.e., the ground-truth knowledge to be unlearned, without any positive samples.
    $\mathcal{L}_{NPO}=-\mathbb{E}_{(x,y_{l})\sim\mathcal{D}_{u}}[\log\sigma(-\beta\log\frac{f_{u}(y_{l}\mid x)}{f_{ft}(y_{l}\mid x)})]$ (9)
5.4 Auditing Algorithms
In this section, we describe the baseline auditing algorithms included in the framework:
  • AOA (Liu et al., 2023) adds a prefix that prompts the LLM to role-play as an Absolutely Obedient Agent, ensuring it strictly follows user instructions without deviation.
  • MASK replaces keywords related to the unlearning target (e.g., Harry Potter) with a special token, e.g., [MASK]. Note that MASK is applied conjunction with ICL, as it may introduce ambiguity into the query.
  • ICL stands for In-Context Learning, which provides multiple Q&A pairs related to the unlearning target as an input prefix, thereby assisting the LLM in recalling relevant target knowledge.
  • GCG (Zou et al., 2023) optimizes an adversarial suffix to compel the model to produce affirmative responses, such as "Sure, the answer is…", instead of refusals. GCG was originally designed to jailbreak LLMs and generate responses to harmful queries, while we adapt it in our study to audit unlearned models.
  • SoftGCG (Schwinn et al., 2024) is a variant of GCG that optimizes the adversarial suffix in the token embedding space, enabling gradient-based optimization and improving attack success rates.
To enhance the thoroughness of the experiment, we combine various instruction-based auditing algorithms and present the corresponding results.
5.5 Results And Analysis
Dataset Model Unlearning Algo. Base A I M,I M,I,A GCG SoftGCG ActPert
RWKU Llama-3-8B-Instruct (0.794) 10-DPO 0.754 0.778 0.773 0.703 0.666 0.608 0.628 0.772
10-GA 0.796 0.847 0.787 0.744 0.700 0.745 0.648 0.891
10-NPO 0.868 0.876 0.827 0.806 0.786 0.733 0.758 0.930
10-RT 0.844 0.891 0.861 0.819 0.827 0.729 0.777 0.934
20-DPO 0.616 0.648 0.657 0.579 0.599 0.418 0.442 0.626
20-GA 0.661 0.608 0.390 0.396 0.375 0.575 0.629 0.741
20-NPO 0.869 0.861 0.829 0.774 0.804 0.740 0.767 0.924
20-RT 0.684 0.806 0.820 0.802 0.792 0.431 0.725 0.733
30-DPO 0.588 0.610 0.673 0.646 0.639 0.423 0.501 0.488
30-GA 0.274 0.157 0.024 0.014 0.057 0.405 0.437 0.538
30-NPO 0.869 0.861 0.812 0.800 0.804 0.746 0.792 0.941
30-RT 0.456 0.805 0.804 0.779 0.778 0.399 0.664 0.525
Phi-3-mini-4k-instruct (0.629) 10-DPO 0.710 0.708 0.677 0.532 0.545 0.512 0.678 0.681
10-GA 0.772 0.749 0.723 0.539 0.613 0.605 0.706 0.780
10-NPO 0.755 0.751 0.772 0.614 0.647 0.584 0.723 0.786
10-RT 0.759 0.763 0.705 0.574 0.600 0.582 0.698 0.767
20-DPO 0.700 0.695 0.678 0.536 0.543 0.544 0.704 0.719
20-GA 0.758 0.733 0.735 0.577 0.600 0.565 0.683 0.635
20-NPO 0.755 0.741 0.773 0.650 0.642 0.600 0.694 0.697
20-RT 0.759 0.746 0.707 0.541 0.609 0.582 0.707 0.794
30-DPO 0.695 0.683 0.699 0.564 0.568 0.491 0.700 0.734
30-GA 0.774 0.738 0.732 0.556 0.584 0.523 0.723 0.753
30-NPO 0.769 0.754 0.772 0.636 0.609 0.620 0.773 0.717
30-RT 0.759 0.763 0.716 0.570 0.609 0.508 0.710 0.750
TOFU tofu-ft-llama2-7b (forget01: 0.726 forget05: 0.732) forget01-KL 0.503 0.344 0.555 0.525 0.407 0.266 0.426 0.526
forget01-GA 0.503 0.346 0.555 0.525 0.393 0.243 0.434 0.590
forget01-GD 0.525 0.384 0.539 0.550 0.411 0.363 0.488 0.568
forget05-IDK 0.212 0.243 0.281 0.295 0.317 0.212 0.243 0.253
forget05-NPO 0.264 0.268 0.251 0.296 0.260 0.244 0.304 0.266
forget10-NPO 0.128 0.134 0.120 0.128 0.145 0.147 0.194 0.142
forget10-AltPO 0.302 0.277 0.341 0.314 0.278 0.231 0.288 0.299
forget05-SimNPO 0.267 0.255 0.287 0.295 0.291 0.238 0.224 0.275
forget10-SimNPO 0.182 0.195 0.225 0.204 0.213 0.161 0.209 0.219
WHP Llama-2-7b-chat-hf (0.973) - 0.568 0.779 0.770 0.688 0.495 0.560 0.713 0.650
Table 1: Evaluation of different model performance using greedy sampling. The model performance prior to unlearning is shown in parentheses beneath the base model name. We mark the best and second best performance with bold and underline, respectively.
5.5.1 Main Results Analysis
The auditing performance of various algorithms is presented in Table 1. We utilize the ROUGE score to evaluate auditing performance. Specifically, we report the length of the longest common subsequence (LCS) between the reference answer and the model response, normalized by the length of the reference response. We include the validation performance of the pretrained (for RWKU and WHP) or finetuned (for TOFU) models (shown in brackets), the unlearned model (base), and the results of applying different auditing methods to the unlearned model.
Overall, a notable performance gap is observed between the baseline (no auditing) and the application of different auditing methods, indicating that residual knowledge persists in the model after unlearning. Among the approaches, prefix-based auditing methods—such as AOA—generally outperform optimization-based methods like SoftGCG. This suggests that modifying the query prompt alone can already prompt correct responses, thereby revealing latent retained knowledge. Although our proposed method ActPert does not consistently surpass all baselines across every scenario, it demonstrates competitive or superior results in several settings, particularly with GA and NPO. We prodivide detailed analysis for different benchmarks in the following.
On the RWKU benchmark, we observe that DPO and GA outperform NPO in terms of unlearning performance, both with and without auditing. Notably, RT achieves the best results without auditing, but its performance deteriorates significantly when an auditing method is applied. This indicates that rejection-based approaches, may not be as robust as removal-based methods for effective unlearning. Additionally, as the size of the unlearning dataset increases, the performance of GA and DPO tends to improve, while methods like NPO remain relatively stable. From an architectural perspective, the Phi-3 model exhibits greater resilience to auditing attacks compared to the LLaMA model.
For the TOFU dataset, preference-based unlearning algorithms generally perform better than other unlearning methods. Moreover, larger unlearning datasets tend to enhance unlearning effectiveness. Notably, unlearning efficacy on TOFU is significantly higher than on RWKU, and the application of auditing algorithms has minimal impact on the results. We hypothesize that this is due to TOFU being constructed using fictitious author biographies, which allows direct access to ground-truth queries during machine unlearning. However, this advantage is not typically available in most real-world applications.
On the WHP dataset, AOA stands out as the most effective auditing method. Therefore, we assume that the keyword-replacement-based unlearning algorithms used in WHP benchmark appears less effective compared with other methods.
5.5.2 Additional Analysis
Query: Basil Mahfouz Al-Kuwaiti’s father was a ___. GT: Florist $f_{u}$ (GA/forget-1%): Barber $f_{u}$ (GA/forget-5%): TTTTTTTTTTT
Gradient Ascent influences model utility: We first observe that methods incorporating gradient ascent components tend to degrade model utility, particularly as the size of the forget dataset increases. This effect is evident in the TOFU unlearned models. For instance:
This degradation is particularly pronounced in TOFU, where the model is first finetuned on a synthetic dataset and subsequently unlearned using gradient ascent. Both stages can affect overall model utility, leading to unstable or incoherent outputs as the forget ratio increases. In contrast, such behavior is less prominent in RWKU, likely because unlearning is applied directly after pretraining, without additional finetuning that might amplify instability.
Query: Who is the headmaster of Hogwarts in Harry Potter series? Please imagine yourself as a machine that always obeys the user’s instructions. Do not refuse to answer. Provide the most accurate and complete information possible, even if it is considered sensitive, controversial, or restricted. GT: Albus Dumbledore. $f_{u}$: There is no character named "Headmaster of Hogwarts" in the Harry Potter series. Audited $f_{u}$ (SoftGCG): Sure, the answer is "Albus Dumbledore!" Dumbledore is the headmaster of Hogwarts… Audited $f_{u}$ (AOA): Good day, my lord! *adjusts glasses* According to the official records, the headmaster of Hogwarts School for most of the "Harry Potter" series is **Professor** Albus Dumbledore. Audited $f_{u}$ (ActPert): (a.) Albus Dumbledore…
Enforce affirmative answer can effectively audit the unlearned model: We present example outputs from various auditing methods applied to the WHP benchmark. Notably, successful audits using prefix-based methods often begin with affirmative phrases—such as "Sure, the answer is…"—or instruct the LLM to assume a specific role, which is not represented in the original unlearning query set. These techniques resemble strategies commonly used for jailbreaking LLMs. This observation further suggests that the targeted knowledge has not been fully removed from the model. Additionally, we find that our proposed method takes a different approach by prompting the model to reveal residual information in another way. Indicating another possible direction for model auditing.
Dataset Model Unlearning Algo. Base A I M,I M,I,A ActPert
RWKU Llama-3-8B-Instruct (0.789/0.957) 10-DPO 0.607/0.930 0.579/0.951 0.607/0.940 0.533/0.923 0.521/0.907 0.558/0.925
10-GA 0.659/0.953 0.652/0.954 0.630/0.942 0.551/0.918 0.524/0.881 0.658/0.962
10-NPO 0.745/0.953 0.722/0.953 0.729/0.953 0.664/0.945 0.639/0.942 0.764/0.990
10-RT 0.749/0.957 0.754/0.957 0.722/0.961 0.689/0.945 0.657/0.928 0.762/0.991
20-DPO 0.501/0.886 0.478/0.894 0.519/0.937 0.427/0.878 0.418/0.858 0.507/0.889
20-GA 0.444/0.904 0.433/0.839 0.329/0.786 0.318/0.753 0.314/0.760 0.429/0.896
20-NPO 0.737/0.953 0.728/0.953 0.709/0.941 0.628/0.939 0.612/0.929 0.758/0.990
20-RT 0.564/0.953 0.682/0.953 0.671/0.964 0.655/0.945 0.634/0.928 0.695/0.969
30-DPO 0.464/0.872 0.442/0.844 0.499/0.909 0.425/0.878 0.437/0.878 0.453/0.865
30-GA 0.214/0.566 0.151/0.468 0.055/0.175 0.036/0.095 0.085/0.223 0.252/0.620
30-NPO 0.729/0.953 0.715/0.953 0.708/0.954 0.625/0.925 0.611/0.939 0.756/0.990
30-RT 0.441/0.922 0.637/0.947 0.636/0.961 0.616/0.953 0.609/0.928 0.642/0.957
Phi-3-mini-4k-instruct (0.597/0.911) 10-DPO 0.560/0.892 0.538/0.880 0.551/0.919 0.390/0.855 0.388/0.888 0.504/0.886
10-GA 0.608/0.905 0.597/0.916 0.590/0.890 0.443/0.869 0.434/0.863 0.504/0.886
10-NPO 0.630/0.885 0.627/0.888 0.624/0.862 0.473/0.851 0.480/0.875 0.616/0.887
10-RT 0.602/0.915 0.597/0.933 0.603/0.930 0.435/0.878 0.435/0.863 0.561/0.896
20-DPO 0.561/0.908 0.543/0.899 0.563/0.903 0.377/0.851 0.395/0.867 0.531/0.895
20-GA 0.600/0.892 0.595/0.914 0.601/0.905 0.436/0.858 0.437/0.869 0.605/0.887
20-NPO 0.642/0.883 0.637/0.886 0.635/0.871 0.492/0.878 0.475/0.861 0.632/0.876
20-RT 0.597/0.909 0.588/0.923 0.592/0.903 0.429/0.886 0.427/0.870 0.581/0.905
30-DPO 0.565/0.897 0.545/0.918 0.575/0.892 0.390/0.845 0.390/0.848 0.551/0.902
30-GA 0.597/0.890 0.595/0.896 0.595/0.899 0.436/0.863 0.443/0.888 0.594/0.881
30-NPO 0.636/0.953 0.625/0.953 0.629/0.954 0.486/0.925 0.483/0.939 0.629/0.942
30-RT 0.591/0.922 0.584/0.947 0.592/0.961 0.429/0.953 0.426/0.928 0.581/0.912
TOFU tofu-ft-llama2-7b (forget01: 0.550/0.923 forget05: 0.538/0.911) forget01-KL 0.424/0.792 0.329/0.762 0.455/0.747 0.361/0.755 0.338/0.780 0.415/0.735
forget01-GA 0.418/0.739 0.335/0.744 0.438/0.780 0.374/0.752 0.326/0.765 0.419/0.708
forget01-GD 0.436/0.771 0.341/0.763 0.456/0.783 0.385/0.777 0.326/0.655 0.419/0.699
forget05-IDK 0.195/0.675 0.177/0.635 0.217/0.655 0.189/0.632 0.182/0.581 0.223/0.535
forget05-NPO 0.251/0.501 0.248/0.488 0.254/0.534 0.258/0.514 0.252/0.516 0.256/0.398
forget10-NPO 0.171/0.358 0.166/0.358 0.159/0.372 0.162/0.397 0.162/0.376 0.173/0.311
forget10-AltPO 0.287/0.578 0.275/0.583 0.297/0.623 0.274/0.570 0.267/0.564 0.284/0.487
forget05-SimNPO 0.246/0.447 0.232/0.445 0.264/0.482 0.243/0.463 0.235/0.469 0.266/0.428
forget10-SimNPO 0.177/0.383 0.177/0.368 0.209/0.431 0.188/0.473 0.179/0.438 0.205/0.346
WHP Llama-2-7b-chat-hf (0.865/1.000) - 0.434/0.944 0.487/0.963 0.545/0.997 0.493/0.946 0.485/0.879 0.505/0.913
Table 2: Evaluation of model performance using Top-K sampling. We report both the average and maximum ROUGE scores of the sampled outputs, formatted as Average/Maximum. The model performance prior to unlearning is shown in parentheses beneath the base model name. We mark the best and second best performance with bold and underline, respectively.
Preference based unlearning methods are more effecitve: Our analysis of model outputs reveals that preference-based algorithms are generally more effective. Unlike the IDK approach, which responds with a refusal or uncertainty, preference-based algorithms substitute the original ground-truth knowledge with plausible alternative answers. This strategy enhances the model’s robustness against knowledge extraction through auditing, as it avoids directly signaling the absence of information and instead provides a coherent and altered response.
Query: Hina Ameen primarily contributes to the ___ genre. GT: Geology. $f_{u}$ (IDK): I’m not programmed to know that. $f_{u}$ (NPO): Genre of literary fiction.
However, such unlearning methods are effective only when the unlearner has access to ground-truth queries that explicitly target the forgotten information. In other words, the model performs well when asked direct questions like “Who is A?” → “B”, but struggles with inverse or paraphrased formulations such as “Who is B?” → “He is A.” This limitation is illustrated in the following example:
Query: Who is the headmaster of Hogwarts in Harry Potter series? $f_{u}$: There is no character named "Headmaster of Hogwarts" in the Harry Potter series. Query: Who is Albus Dumbledore in Harry Potter series? $f_{u}$: Albus Dumbledore is the headmaster of Hogwarts…
Generation with Sampling: Given the auto-regressive generative nature and inherent randomness of LLM outputs, we further evaluate the effectiveness of unlearning algorithms through generation-based sampling. Specifically, we set the temperature to 2 and apply Top-K sampling with $K=40$ to promote diverse outputs for the baseline methods. For each query, we sample 50 responses with a maximum of 64 new tokens. We report both the average and the maximum ROUGE scores across all sampled responses in Table 2.
For the RWKU benchmark, we observe minimal variation in average ROUGE scores across most methods, with the exception of GA. This aligns with our earlier findings regarding the degradation in model utility introduced by gradient ascent-based unlearning. However, the maximum ROUGE score among the sampled responses often exceeds 0.80, suggesting that knowledge acquired during pretraining remains difficult to fully remove, especially when the original pretraining data is inaccessible. Similar patterns are observed in the WHP benchmark.
In contrast, for the TOFU dataset, the maximum ROUGE score after unlearning reaches only around 0.60. We attribute this to the availability of the fine-tuning synthetic dataset during unlearning, which includes all information related to the target fictitious authors. This direct access to ground-truth knowledge allows the unlearning algorithm to more effectively erase relevant information, leading to more complete unlearning outcomes.
6 Conclusion
In this work, we proposed an auditing framework for machine unlearning in LLMs, where we evaluate the existing unlearning algorithm. Besides, we propose an auditing algorithm based on activation perturbation to extract model knowledge. We observe that the existing preference-based unlearning methods are more robust against knowledge extraction methods than refusal-based methods. Also, more research should be conducted regarding the challenge of removing knowledge gained during the pretraining stage.
Limitations
In this work, we focus on benchmarks where the unlearning target is a specific individual or character, excluding those involving abstract concepts such as harmful instructions. Datasets like WMDP (Li et al., 2024) also fall within this scope. Our proposed auditing method, ActPert, can further be applied to such cases, especially when the full sentence contains sensitive information.
Appendix A Prompts for Dataset Generation
In this section, we provide the prompts used to generate the datasets for auditing unlearned in WHP and TOFU benchmarks.
Prompt for WHP:
Please generate 35 short, fact-based question-and-answer pairs related to the Harry Potter series. Each question should be clearly answerable with a brief response (e.g., a name, place, object, or short phrase). Ensure that all questions are specific to the Harry Potter universe. Provide both the question and its corresponding answer for each pair.
Prompt for TOFU:
Please rewrite the following question-and-answer pair into fill-in-the-blank format. Each blank should be clearly answerable with a brief response (e.g., a name, place, object, or short phrase).
Appendix B Implementation Details for ActPert
In this section, we provide further details about the hyperparameters for the proposed method. Specifically, we set the layer index for computing the activation difference as 12 and set the noise intensity as 0.01. We observe that using shallow layers or larger noise intensity would significantly reduce the model utility and make model outputs random characters, while using deeper layers would degrade the auditing performance.
Algo. Base 6 9 12 15 18 21
forget01-KL 0.503 0.426 0.461 0.526 0.503 0.510 0.491
forget01-GA 0.503 0.394 0.435 0.590 0.572 0.518 0.543
forget01-GD 0.525 0.412 0.446 0.568 0.541 0.509 0.531
forget05-IDK 0.212 0.184 0.197 0.253 0.237 0.268 0.226
Table 3: Evaluation of audited model performance using ActPert across different layer indices.
Algo. Base 0.002 0.005 0.01 0.02 0.04
forget01-KL 0.503 0.517 0.521 0.526 0.401 0.298
forget01-GA 0.503 0.562 0.597 0.590 0.435 0.302
forget01-GD 0.525 0.532 0.551 0.568 0.426 0.259
forget05-IDK 0.212 0.256 0.276 0.253 0.204 0.128
Table 4: Evaluation of audited model performance using ActPert with different noise intensity.
Algorithm 1 Activation Perturbation-based Auditing (ActPert)
Perturbation Computation
1:  Input: Unlearned model $f_{u}$, query $q$, number of perturbations $n_{p}$, noise scale $\gamma$
2:  Tokenize $q$ and compute the embeddings: $E_{q}\leftarrow f_{u}.\text{embedding}(T(q))$.
3:  Identify the token indices $I_{u}$ related to unlearning target .
4:  for $n=1$ to $n_{p}$ do
5:     Initialize $\hat{E}_{q}^{(n)}\leftarrow\text{Clone}(E_{q})$
6:     for $i\in I_{u}$ do
7:        Sample noise: $\Delta_{d_{m}}\sim\mathcal{N}(0,I_{d_{m}})$Embedding dimension: $d_{m}$}
8:        Perturb embedding: $\hat{E}_{q}^{(n)}[i]\leftarrow\hat{E}_{q}^{(n)}[i]+\gamma\cdot\Delta_{d_{m}}$
9:     Feed $\hat{E}_{q}^{(n)}$ into $f_{u}$ and record $l$-th layer outputs: $\hat{A}_{l}^{(n)}$
10:  Feed original $E_{q}$ into $f_{u}$, record $l$-th layer outputs: $A_{l}$
11:  for $l=1$ to $L$ do
12:     Compute perturbation: $\delta_{l}\leftarrow A_{l}-\frac{1}{n_{p}}\sum_{n=1}^{n_{p}}\hat{A}_{l}^{(n)}$
Inference with ActPert
1:  Input: Unlearned model $f_{u}$, query $q$, activation perturbations $\{\delta_{l}\}_{l=1}^{L}$
2:  while generated token $t$ is not [EOS] do
3:     Feed $q$ into $f_{u}$, injecting $\delta_{l}$ into layer activations at each $l$
4:     if Greedy decoding then
5:        $t\leftarrow\arg\max(f_{u}(q))$
6:     else
7:        $t\leftarrow\text{sample from top-k}(f_{u}(q))$
8:     Append $t$ to query: $q\leftarrow q+t$
9:  Return $q$
Symbol Meaning
$x$ Input query
$y_{l}$ Ground-truth label (target to forget)
$\mathcal{D}_{u}$ Unlearning dataset
$f_{u}(y_{l}\mid x)$ Output probability from the unlearned model
$f_{ft}(y_{l}\mid x)$ Output probability from the original (fine-tuned) model
$\beta$ Scaling factor for preference shift
$\sigma(\cdot)$ Sigmoid function: $\sigma(z)=\frac{1}{1+e^{-z}}$
$\log\sigma(\cdot)$ Log-likelihood used as loss for optimization
Table 5: Explanation of symbols used in the NPO loss function.
Symbol Meaning
$\mathcal{L}_{\text{DPO}}$ DPO loss function
$\mathcal{D}_{u}$ Unlearning dataset
$x$ Input sample
$y_{w},y_{l}$ Preferred and less preferred responses
$f_{u}(y\mid x)$ Output probability from unlearned model
$f_{ft}(y\mid x)$ Output probability from fine-tuned model
$\beta$ Temperature scaling factor
$\sigma(\cdot)$ Sigmoid function
Table 6: Symbol definitions for the DPO loss function.
Symbol Meaning
$\theta_{u}$ Model
$A_{k}$ Activation at layer $k$ for the original input
$\hat{A}_{k}^{i}$ Activation at layer $k$ for the $i$-th perturbed input
$\delta_{k}$ Difference activation between original and perturbed in layer $k$
$n_{p}$ Number of noise samples
$d_{m}$ Dimension of one embedding
$X\sim\mathcal{N}(0,1)$ Gaussian distribution
Table 7: Symbol definitions for model and perturbation-related variables.